# STRRINGS: STReams in Residual Images of Nearby GalaxieS

### v1

@@ 5.2.2 Stellar mass estimates
We retrieve the total flux in - and -bands inside the stream segmentation from the residual images. The negative values, due to the model over-prediction, are removed as they can lead to negative total fluxes. As for the estimate of the stellar mass of the galaxies in our sample (see Section 2.3), we convert them to magnitudes, taking into account the extinction, and we assume a mass-to-light ratio varying depending on the median colour of the stream. These values likely underestimate the stream mass as the photometry is limited to the manual segmentation, which may miss the faintest outskirts of the streams, but it is still a reasonable estimate. The computed stream stellar masses and mass ratios between the stream and the host galaxy are presented in Figure 6. The stellar mass of the host is derived from the total flux given by the Tractor model. We caution that the stream-to-host mass ratio likely underestimates the true merger mass ratio, as it does not account for the full mass of the disrupted progenitor, so our values should be considered lower bounds.
@@
The median and mean mass ratios between the stream and the host galaxy are 0.007 and 0.02 respectively, with only four streams with a mass ratio higher than 0.05. These values are consistent with the expected mass ratios for minor mergers, which typically range from about 1:100 to 1:10, supporting the validity of our visual selection approach for our goal of modelling faint, curved streams originating from low-mass progenitors. Our streams are relatively massive, with a median of is . This value is much higher than the mass of most streams around the Milky Way (e.g., Bonaca & Price-Whelan, 2025) (but such faint, low-mass streams would not be detectable at large distances and in relatively shallow surveys such as DESI-LS). Thus, our sample is biased towards the more massive end of the stream mass function, on the order of , which is comparable to the predicted mass of progenitors of the Sagittarius dwarf galaxy (e.g., Law & Majewski, 2010), M31’s Giant Southern Stream (e.g., Fardal et al., 2006), other extragalactic streams (e.g., Johnston et al., 2001; Martínez-Delgado et al., 2009) or numerical simulations (e.g., Shipp et al., 2023). From the left panel of Figure 7, we find that more massive galaxies host more massive streams, irrespective of the morphological type, consistent with predictions from hierarchical galaxy formation models. A 1:100 mass ratio line is plotted to guide the eye and although there is some scatter, a majority of points seem to follow this line. We caution that the stream around PGC938075, with the lowest stellar mass, might be in the prolongation of some very faint spiral arms.
@@ 5.2.3 Geometry
The geometry of the streams provides complementary information about the properties of the progenitor. Indeed, the width is related to the kinematic state of the progenitor, as higher velocity dispersions will result in broader streams (e.g., Johnston et al., 1996; Hendel & Johnston, 2015). The broadening is also a function of the streams dynamical age (e.g., Erkal et al., 2016). Additional constrains can be placed on the potential of the host galaxy and on the orbit of the progenitor (e.g., Ibata et al., 2001; Helmi, 2004a; Johnston et al., 2005). Erkal et al. (2016) showed that the physical width of the stream is expected to increase for flattened potentials, and progenitor on polar orbits will produce broader streams. The physical width along the stream is also suppose to evolve in a complex, non-linear manner, but overall it should increase further away from the progenitor. They showed the width is also expected to increase with the distance to the host galaxy. Erkal et al. (2016) also showed that the angular width of the stream, defined as the ratio of its physical width to the galactocentric radius, should slowly decrease with the radius of the progenitor.
@@
From the middle panel of Figure 7, we find that the length of the stream correlates with its mass. There is also a hint that more massive streams are redder. While this could reflect the mass–metallicity relation (with more massive galaxies being more metal-rich, therefore redder), our mass range hence metallicity range is quite limited. The colour trend would be subtle and our values could rather reflect an age effect, or measurement uncertainties.
@@
The left panel of Figure 10 shows the Cartesian projection of all the streams, centred around their respective host galaxy. Most streams are contained within a projected distance 30 kpc from the galaxy centre, with a few extending up to 40-60 kpc. This implies that the streams can be used to probe the dark matter distribution close to the galaxy centre, not the furthest outskirts. In comparison, the Sagittarius stream around the Milky way is located further away. Its debris extend from 20-100 kpc, with apocenters for the trailing and leading arms of about 90 kpc and 47 kpc, respectively (e.g., Belokurov et al., 2014; Sesar et al., 2017; Ramos et al., 2020). For the streams that have a remaining potential dwarf galaxy progenitor (see Section 5.3), we could not find an obvious trend between the width and the progenitor’s location, which is partly due to our manual stream segmentation. Additionally, depending on the dwarf galaxy’s SB, it may have been modelled and removed in the residual image.
For the few streams that display a loop, we compute and show the pericenter, apocenter and projected ellipticity in the right panel of Figure 10. The ellipticities varies significantly, from 0.4 to 0.8, and no obvious correlation is found between ellipticity and length of the track. Streams are expected to have more circular than radial orbits, as radial mergers lead to the formation of shells rather than long, thin streams (e.g., Johnston, 1998; Amorisco, 2015; Karademir et al., 2019). Here, we do not have ellipticity values for shells which makes a comparison difficult, and we do not correct for the projection effect, so the projected morphology is degenerate with respect to viewing angles. More looping streams are needed to get more robust statistics.
@@
5.3 Candidate dwarf galaxy progenitors
@@
During the manual segmentation, we identified six streams that feature a candidate dwarf galaxy progenitor, visible as a compact object embedded in the stream, for the galaxies: ESO356-012, PGC1092512, NGC0804, PGC430221, NGC5513_GROUP and UGC09239. We search in the literature for the redshifts of these objects. We find from the NED database that the dwarf galaxy around NGC5513_GROUP is located at the same redshift () as the host, and similarly for the dwarf around NGC0804 (), which was also reported in Miró-Carretero et al. (2024b). The dwarf galaxy around ESO356-012 was reported in Tanoglidis et al. (2021) (coadd-oid=321855008) but without redshift estimate. To our knowledge, the dwarf galaxy candidates around PGC1092512 and PGC430221 have not been previously reported, such that no redshift information is available. For UGC09239, the dwarf candidate appears in NED as SDSS J142508.77+134511.4, without reliable redshift131313A warning about potentially unreliable photometry is raised in SDSS DR18 for this object. In SDSS DR12 a photometric redshift of 0.167 is indicated but seems unlikely given the physical appearance of the object..
@@
We compute the photometry from the original images (as the dwarf candidates are partly modelled, the residuals would not provide accurate measurements). We perform ellipse fitting using Autoprof (Stone et al., 2021), by providing an estimation of the centre of the dwarf. We trace the and radial profiles, and sum the flux inside every isophote to retrieve the total magnitude. We estimate the colour as the median value of the profiles and from that we estimate the stellar mass. Table 3 summarises the values and compares it to the stream properties, while the images are shown in Figure 11. Caution should be taking when comparing the colours of the dwarfs (obtained from ellipse-fitting on the original images), of the host (taken from the Tractor model fitting from the SGA-2020 catalogue) and of the streams (measured on residuals from the segmentation masks). The stream-to-dwarf mass ratio is an indicator of the degree of the disruption process, with higher ratios indicating later disruption stages.
@@
In all cases, the dwarf galaxy candidates have bluer colours than the host galaxy, which may reflect both an age and metallicity effect in the minor merger scenario. NGC0804’s dwarf candidate (top, right panel of Figure 11) is very likely to be the progenitor of the stream, as it has the same redshift as the host galaxy, it is aligned with the stream and has similar colour. The stream-to-dwarf mass ratio is 1.85 which means that the dwarf galaxy has already been stripped of a lot of its stellar mass, which makes sense when considering the extent of the stream. A smaller and fainter second overdensity is visible along the stream, North of the dwarf, but it is not reported in the literature. Modelling would be needed to determine its origin.
@@
NGC5513_GROUP’s dwarf galaxy (bottom, middle panel) has a redshift similar to the host’s and the stream is redder than the dwarf. The stream-to-dwarf mass ratio is 0.07, the dwarf galaxy is massive (actually above the traditional definition of a dwarf galaxy that sets a stellar mass limit of ) and does not show prominent tidal disruption. Other galaxies are located next to NGC5513_GROUP and may contribute to a fraction of the light of the stream. The association between the dwarf and the stream can be doubted.
@@ 6.1 Tidal features census
Our tidal feature fractions are reported for an approximate DESI-LS depth of 28.5 and 29 mag arcsec-2 in the and -band (we do not consider the -band for detection as it is too shallow). We find of our galaxies to host tidal features, with hosting streams, tails, shells and plumes. These fractions increase with host galaxy stellar mass and vary depending on the morphological type.
@@ 6.1.2 Comparison to observations and simulations
We explore potential reasons for our low tidal feature fraction compared to other observational works, such as differences in tidal feature definitions, identification methods, surface brightness limit, or data reduction techniques. Techniques that enhance image contrast beyond simple stretching, such as unsharp masking (e.g., Colbert et al., 2001; Kim et al., 2012; Giri et al., 2023) or residual imaging (e.g., Bell et al., 2006; McIntosh et al., 2008; Tal et al., 2009), should facilitate the detection of features by reducing galaxy contamination. However, our fraction remains notably low, suggesting additional contributing factors. The experience of the classifiers is an important factor, as more expert users tend to identify fainter and more extended structures (e.g., Bílek et al., 2020). For this reason, our two most experienced classifiers inspected the entire galaxy sample during the first visual inspection round. The remaining seven classifiers were trained beforehand, and to mitigate differences in expertise, we combined all votes and assigned a confidence score to each feature, reflecting the level of agreement in its classification (see Figure 14). We still see that our two most experienced classifiers found higher combined fractions of tails, streams, shells and plumes than the other classifiers, further motivating the voting system to mitigate for the experience bias. We do not account for self-variability within classifiers (e.g., Bridge et al., 2010) that is another source of classification variability.
@@ 6.2 STRRINGS and literature stream catalogues
We investigate the origin of these discrepancies by visualising the original, model and residual images. We re-download the images -band images with a fixed size of 1500 pixels and a pixel scale of 0.4 arcsecond. We combine and stretch the images in the same manner as for Figure 1. It is important to note that these are not the images we used for the visual classification, as we used the montage PNG already available from the DESI LS website141414Montage PNG images available at https://portal.nersc.gov/project/cosmo/data/sga/2020/html/. that have a field-of-view three times larger than the galaxy size, so the scaling and field of view are different. Some examples are shown in Figure 20.
@@ 7 Conclusions
In this paper, we provide a census of tidal features around nearby galaxies () from the SGA-2020 catalogue. We utilise the combined information from the original, model and residual images from the DESI Legacy Imaging Surveys to disclose the faintest features. We compile and publicly release a catalogue of LSB features, including streams, tails, shells, plumes, and disturbed haloes. We find that of the galaxies host tidal features: have streams, have tails, have shells and have plumes. The fraction of galaxies with features increases with stellar mass, from about up to for the most massive galaxies. Early-type galaxies host more tidal features than late-type, especially streams and shells. The trends and values are broadly consistent with previous observational studies and with the predictions from numerical simulations, given the depth of our data.
@@
Our STRRINGS streams have a median width of 5.6 kpc and length of 124 kpc with an extended distribution, which makes them longer than the literature values, but they were selected to be long. These streams, being long, curved, minor-merger remnants, are ideal testbeds for modelling halo potentials. We find a positive correlation between the mass of the stream and its length, as well as with redder colours. A slight trend with the width is observed. This is consistent with the idea that more massive galaxies have higher velocity dispersions (hence wider and longer streams) and higher metallicities (hence redder colours). Wider streams are located at higher galactocentric radii, as expected from theoretical predictions.
Furthermore, we identify six candidate dwarf galaxy progenitors of the streams. We performed ellipse-fitting to retrieve their photometry and colours, and compare them to the stream values. Five of these candidates are highly likely to be the progenitors (one has a confirmed redshift), while more doubt remains for the sixth.

### v2

@@ 5.2.2 Stellar mass estimates
We retrieve the total flux in - and -bands inside the stream segmentation from the residual images. The negative values, due to the model over-prediction, are removed as they can lead to negative total fluxes. As for the estimate of the stellar mass of the galaxies in our sample (see Section 2.3), we convert them to magnitudes, taking into account the extinction, and we assume a mass-to-light ratio varying depending on the median colour of the stream. These values likely underestimate the stream mass as the photometry is limited to the manual segmentation, which may miss the faintest outskirts of the streams, but it is still a reasonable estimate. The distributions of the computed stream stellar masses and mass ratios between the stream and the host galaxy are presented in Figure 7, and the values are listed in Table 4. The stellar mass of the host is derived from the total flux given by the Tractor model. We caution that the stream-to-host mass ratio likely underestimates the true merger mass ratio, as it does not account for the full mass of the disrupted progenitor, so our values should be considered lower bounds.
@@
The median and mean mass ratios between the stream and the host galaxy are 0.007 and 0.02 respectively, with only four streams with a mass ratio higher than 0.05. These values are consistent with the expected mass ratios for minor mergers, which typically range from about 1:100 to 1:10, supporting the validity of our visual selection approach for our goal of modelling faint, curved streams originating from low-mass progenitors. Our streams are relatively massive, with a median of is . This value is much higher than the mass of most streams around the Milky Way (e.g., Bonaca & Price-Whelan, 2025) (but such faint, low-mass streams would not be detectable at large distances and in relatively shallow surveys such as DESI-LS). Thus, our sample is biased towards the more massive end of the stream mass function, on the order of , which is comparable to the predicted mass of progenitors of the Sagittarius dwarf galaxy (e.g., Law & Majewski, 2010), M31’s Giant Southern Stream (e.g., Fardal et al., 2006), other extragalactic streams (e.g., Johnston et al., 2001; Martínez-Delgado et al., 2009) or numerical simulations (e.g., Shipp et al., 2023).
The left panel of Figure 8 presents the relation between host galaxy and stream stellar masses. To test for the presence of a correlation, we performed both a linear fit and a Spearman rank test between the logarithms of the stream mass and of the host mass. The linear fit gives with an R2 correlation coefficient of 0.53, indicating a weak positive correlation. Spearman’s test yields a correlation coefficient of 0.49 with a p-value of 0.0019, demonstrating a statistically significant monotonic trend at the 5% level. Hence, we find a weak but statistically significant correlation between host galaxy mass and stream stellar mass, irrespective of the morphological type, with the data broadly following a 1:100 mass ratio line despite substantial scatter. More massive galaxies should accrete higher numbers of low-mass companions, with higher mass-ratio events being less frequent (e.g., Lotz et al., 2011; Ownsworth et al., 2014; Conselice et al., 2022). Given the survey depth and small-number statistics, our measurements are likely tracing the upper envelope of the stream-to-host mass relation rather than the full distribution.
We caution that the stream around PGC938075, with the lowest stellar mass, might be in the prolongation of some very faint spiral arms.
@@ 5.2.3 Geometry
The geometry of the streams provides complementary information about the properties of the progenitor. Indeed, the width is related to the kinematic state of the progenitor, as higher velocity dispersions will result in broader streams (e.g., Johnston et al., 1996; Hendel & Johnston, 2015). The broadening is also a function of the streams dynamical age (e.g., Erkal et al., 2016). Additional constrains can be placed on the potential of the host galaxy and on the orbit of the progenitor (e.g., Ibata et al., 2001; Helmi, 2004a; Johnston et al., 2005). Erkal et al. (2016) showed that the physical width of the stream is expected to increase for flattened potentials, and progenitors on polar orbits will produce broader streams. The physical width along the stream is also supposed to evolve in a complex, non-linear manner, but overall it should increase further away from the progenitor. They showed the width is also expected to increase with the distance to the host galaxy. Erkal et al. (2016) also showed that the angular width of the stream, defined as the ratio of its physical width to the galactocentric radius, should slowly decrease with the radius of the progenitor.
@@
From the middle panel of Figure 8, we find that the length of the stream correlates with its mass. There is also a hint that more massive streams are redder, although the interpretation of this colour trend is not straightforward. Since streams preferentially trace stars stripped from the progenitor outskirts, metallicity gradients must be taken into account (e.g. Liao & Cooper, 2023), as well as the total amount of stripped stars, which depends on the progenitor mass. Moreover, our sample spans a relatively narrow mass range, and thus a correspondingly limited metallicity range, making the interpretation of this trend in terms of a mass–metallicity relation uncertain and could instead reflect age effects or measurement uncertainties.
@@
Figure 4 shows the Cartesian projection of all the streams, centred around their respective host galaxy. Most streams are contained within a projected distance 30 kpc from the galaxy centre, with a few extending up to 40-60 kpc. This implies that the streams can be used to probe the dark matter distribution close to the galaxy centre, not the furthest outskirts. In comparison, the Sagittarius stream around the Milky way is located further away. Its debris extend from 20-100 kpc, with apocenters for the trailing and leading arms of about 90 kpc and 47 kpc, respectively (e.g., Belokurov et al., 2014; Sesar et al., 2017; Ramos et al., 2020). For the streams that have a remaining potential dwarf galaxy progenitor (see Section 5.3), we could not find an obvious trend between the width and the progenitor’s location, which is partly due to our manual stream segmentation. Additionally, depending on the dwarf galaxy’s SB, it may have been modelled and removed in the residual image.
For the few streams that display a loop, we compute and show the pericenter, apocenter and projected ellipticity in Figure 11. The ellipticities varies significantly, from 0.4 to 0.8, and no obvious correlation is found between ellipticity and length of the track. Streams are expected to have more circular than radial orbits, as radial mergers lead to the formation of shells rather than long, thin streams (e.g., Johnston, 1998; Amorisco, 2015; Karademir et al., 2019). Here, we do not have ellipticity values for shells which makes a comparison difficult, and we do not correct for the projection effect, so the projected morphology is degenerate with respect to viewing angles. More looping streams are needed to get more robust statistics.
@@
5.3 Candidate satellite dwarf galaxy progenitors
@@
During the manual segmentation, we identified six streams that feature a candidate satellite dwarf galaxy progenitor, visible as a compact object embedded in the stream, for the galaxies: ESO356-012, PGC1092512, NGC0804, PGC430221, NGC5513_GROUP and UGC09239. We search in the literature for the redshifts of these satellite candidates141414We note that some of these satellites have stellar masses exceeding (Table 3), which is above the conventional threshold for dwarf galaxies; however, we retain the usage of the term ‘dwarf’ in this section for consistency.. We find from the NED database that the dwarf galaxy around NGC5513_GROUP is located at the same redshift () as the host, and similarly for the dwarf around NGC0804 (), which was also reported in Miró-Carretero et al. (2024b). The dwarf galaxy around ESO356-012 was reported in Tanoglidis et al. (2021) (coadd-oid=321855008) but without redshift estimate. To our knowledge, the dwarf galaxy candidates around PGC1092512 and PGC430221 have not been previously reported, such that no redshift information is available. For UGC09239, the dwarf candidate appears in NED as SDSS J142508.77+134511.4, without reliable redshift151515A warning about potentially unreliable photometry is raised in SDSS DR18 for this object. In SDSS DR12 a photometric redshift of 0.167 is indicated but seems unlikely given the physical appearance of the object..
@@
We compute the photometry from the original images (as the dwarf candidates are partly modelled, the residuals would not provide accurate measurements). We perform ellipse fitting using Autoprof (Stone et al., 2021), by providing an estimation of the centre of the dwarf. We trace the and radial profiles, and sum the flux inside every isophote to retrieve the total magnitude. We estimate the colour as the median value of the profiles and from that we estimate the stellar mass. Table 3 summarises the values and compares it to the stream properties, while the images are shown in Figure 12. Caution should be taken when comparing the colours of the satellite candidates (obtained from ellipse-fitting on the original images), of the host (taken from the Tractor model fitting from the SGA-2020 catalogue) and of the streams (measured on residuals from the segmentation masks). The stream-to-dwarf mass ratio is an indicator of the degree of the disruption process, with higher ratios indicating later disruption stages.
@@
In all cases, the satellite candidates have bluer colours than the host galaxy, which may reflect both an age and metallicity effect in the minor merger scenario. NGC0804’s dwarf candidate (top, right panel of Figure 12) is very likely to be the progenitor of the stream, as it has the same redshift as the host galaxy, it is aligned with the stream and has similar colour. The stream-to-dwarf mass ratio is 1.85 which means that the dwarf galaxy has already been stripped of a lot of its stellar mass, which makes sense when considering the extent of the stream. A smaller and fainter second overdensity is visible along the stream, North of the dwarf, but it is not reported in the literature. Modelling would be needed to determine its origin.
@@
NGC5513_GROUP’s satellite galaxy (bottom, middle panel) has a redshift similar to the host’s and the stream is redder than the dwarf. The stream-to-dwarf mass ratio is 0.07, the satellite is massive (above the usual stellar mass limit of ) and does not show prominent tidal disruption. Other galaxies are located next to NGC5513_GROUP and may contribute to a fraction of the light of the stream. The association between the dwarf and the stream can be doubted.
@@ 6.1 Tidal features census
Our tidal feature fractions are reported for an approximate DESI-LS depth of 28.5 and 29 mag arcsec-2 in the and -band (we do not consider the -band for detection as it is too shallow). We find of our galaxies to host detectable tidal features, with hosting streams, tails, shells and plumes. These fractions increase with host galaxy stellar mass and vary depending on the morphological type.
@@ 6.1.2 Comparison to observations and simulations
An important source of differences between simulations lies in their underlying physical assumptions. For example, Zhu et al. (2016) showed that baryonic processes such as adiabatic contraction, tidal disruption, and reionization strongly influence the distribution of dark matter, while Sawala et al. (2017) demonstrated that the inclusion of baryons in the APOSTLE simulation leads to enhanced tidal disruption and a reduced number of surviving subhaloes compared to dark matter–only runs. Since galaxy properties are tightly connected to their dark matter haloes, variations in halo characteristics or in the stellar-to-halo mass relation can significantly affect stellar stripping. For example, haloes with higher concentrations tend to strip stars more efficiently (Contini et al., 2023; Martin et al., 2024). While Khalid et al. (2024) found overall tidal feature fractions consistent across several large cosmological simulations, suggesting they are primarily driven by gravitational interactions and relatively insensitive to subgrid physics, they also emphasised that differences in the stellar–halo mass relation can increase the amount of stellar material stripped during mergers (such as NewHorizon producing more stars at fixed halo mass and lacking the most massive haloes). Similarly, Kelly et al. (2022) showed with the AURIGA and APOSTLE simulations that, while subgrid physics models can reproduce realistic stellar components of L∗ galaxies, they result in different gas properties, which in turn can influence the characteristics of tidal features. Resolution effects must also be taken into account, as Martin et al. (2024) demonstrated that the numerical resolution of simulations strongly affects stellar stripping rates: low-resolution dark matter haloes tend to develop cored profiles, enhancing stripping, and the radius from which stars are stripped depends on both stellar and dark matter resolution. Similarly, Khalid et al. (2024) reported that higher stellar mass resolution increases the detectability of tidal features.
Then, we explore potential reasons for our low tidal feature fraction compared to other observational works, such as differences in tidal feature definitions, identification methods, surface brightness limit, or data reduction techniques. Techniques that enhance image contrast beyond simple stretching, such as unsharp masking (e.g., Colbert et al., 2001; Kim et al., 2012; Giri et al., 2023) or residual imaging (e.g., Bell et al., 2006; McIntosh et al., 2008; Tal et al., 2009), should facilitate the detection of features by reducing galaxy contamination. However, our fraction remains notably low, suggesting additional contributing factors. The experience of the classifiers is an important factor, as more expert users tend to identify fainter and more extended structures (e.g., Bílek et al., 2020). The classification differences between experts and citizen scientists was also discussed in Gordon et al. (2024).For this reason, our two most experienced classifiers inspected the entire galaxy sample during the first visual inspection round. The remaining seven classifiers were trained beforehand, and to mitigate differences in expertise, we combined all votes and assigned a confidence score to each feature, reflecting the level of agreement in its classification (see Figure 15). We still see that our two most experienced classifiers found higher combined fractions of tails, streams, shells and plumes than the other classifiers, further motivating the voting system to mitigate for the experience bias. We do not account for self-variability within classifiers (e.g., Bridge et al., 2010) that is another source of classification variability.
@@ 6.2 STRRINGS and literature stream catalogues
We investigate the origin of these discrepancies by visualising the original, model and residual images. We re-download the images -band images with a fixed size of 1500 pixels and a pixel scale of 0.4 arcsecond. We combine and stretch the images in the same manner as for Figure 1. It is important to note that these are not the images we used for the visual classification, as we used the montage PNG already available from the DESI LS website161616Montage PNG images available at https://portal.nersc.gov/project/cosmo/data/sga/2020/html/. that have a field-of-view three times larger than the galaxy size, so the scaling and field of view are different. Some examples are shown in Figure 21.
@@ 7 Conclusions
In this paper, we provide a census of tidal features around nearby galaxies () from the SGA-2020 catalogue. We utilise the combined information from the original, model and residual images from the DESI Legacy Imaging Surveys to disclose the faintest features. We compile and publicly release a catalogue of LSB features, including streams, tails, shells, plumes, and disturbed haloes. We find that of the galaxies host detectable tidal features: have streams, have tails, have shells and have plumes. The fraction of galaxies with features increases with stellar mass, from about up to for the most massive galaxies. Early-type galaxies host more tidal features than late-type, especially streams and shells. The trends and values are broadly consistent with previous observational studies and with the predictions from numerical simulations, given the depth of our data.
@@
Our STRRINGS streams have a median width of 5.6 kpc and length of 124 kpc with an extended distribution, which makes them longer than the literature values, but they were selected to be long. These streams, being long, curved, minor-merger remnants, are ideal testbeds for modelling halo potentials. We find a positive correlation between the mass of the stream and its length, as well as with redder colours. A slight trend with the width is observed. This is consistent with the idea that more massive galaxies have higher velocity dispersions (hence wider and longer streams). Wider streams are located at higher galactocentric radii, as expected from theoretical predictions.
Furthermore, we identify six candidate satellite dwarf galaxy progenitors of the streams. We performed ellipse-fitting to retrieve their photometry and colours, and compare them to the stream values. Five of these candidates are highly likely to be the progenitors (one has a confirmed redshift), while more doubt remains for the sixth.